OpenStudy (donnie1999):

------------WILL FAN AND MEDAL------------ A mail distribution center processes as many as 180,000 pieces of mail each day. The mail is sent via ground and air. Each land carrier takes 2000 pieces per load. The loading equipment is able to handle as many as 150 loads per day. Let x be the number of loads by land carriers and y the number of loads by carriers. Which system of inequalities represents this situation?

OpenStudy (anonymous):

Any answer choices?

OpenStudy (donnie1999):

\[A. 1250+2000y \ge 180,000. x+y \ge 150\]

OpenStudy (anonymous):

are any of them 2000x+1250y<=180000 x+y<=150?